OpenStudy (anonymous):

Determine if (1/2.-4) is solution to 4x-3y=-10. Please Help... Show work

OpenStudy (anonymous):

substitute the values of x and y in the given equation...... LHS = 4(1/2) - 3(-4) = 2 + 12 = 14 which is not equal to RHS i.e. -10 so (1/2, -4) is not a solution to the given equation